OpenStudy (gabylovesyou):

What is the simplified form of the seventh root of x times the seventh root of x times the seventh root of x times the seventh root of x? 4 times the seventh root of x x to the four-sevenths power the seventh root of 4x x to the seven fourths power

OpenStudy (gabylovesyou):

@sleepyjess

OpenStudy (sleepyjess):

\(^7\sqrt x *^7\sqrt x*^7\sqrt x*^7\sqrt x\)?

OpenStudy (superman499):

what do u think

OpenStudy (gabylovesyou):

i think its B

OpenStudy (superman499):

well u need to go with wat u think it is its ok to do an educated guess some time

OpenStudy (anonymous):

im thinking B to ^-^